OpenStudy (anonymous):

lim(x→∞)⁡(1-2x^2)/((4x+3)^2) Thank you Solomon, you are awesome. I have one more I need help with. Thank you again

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{v \rightarrow ~\infty}\frac{1-2x^2}{(4x+3)^2}}\) this ?

OpenStudy (solomonzelman):

I mean, \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~\infty}\frac{1-2x^2}{(4x+3)^2}}\) . Like this ?

OpenStudy (solomonzelman):

I can't help you, if you don't even let me know if my interpretation is correct or not.

OpenStudy (anonymous):

yes you are correct with the second entry, sorry had to take dog out

OpenStudy (anonymous):

Expand the denominator by multiplying it out

OpenStudy (anonymous):

that's about all I know how to do lol, just started this today so I am not to good at it yet

OpenStudy (anonymous):

Okay, did you expand it?

OpenStudy (anonymous):

16x^2+9

OpenStudy (anonymous):

That isn't correct

OpenStudy (solomonzelman):

you forgot the middle term, @ingallina

OpenStudy (anonymous):

16x^2+24+9

OpenStudy (solomonzelman):

Ok, I thought of... \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~\infty}\frac{1-2x^2}{(4x+3)^2}}\) \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~\infty}\frac{-(2x^2-1)}{(4x+3)^2}}\) \(\large\color{slate}{-\left[ \displaystyle\lim_{x \rightarrow ~\infty}\frac{(2x^2-1)}{(4x+3)^2}\right]}\) then L'H'S twice

OpenStudy (anonymous):

what does lhs twice mean? and I do not understand why you have same problem three different ways? I just need to solve it showing steps to help me understand

OpenStudy (solomonzelman):

I mean to differentiate top and bottom of the limit twice

OpenStudy (solomonzelman):

Do you understand how I got, \(\large\color{slate}{-\left[ \displaystyle\lim_{x \rightarrow ~\infty}\frac{(2x^2-1)}{(4x+3)^2}\right] }\)

OpenStudy (solomonzelman):

??

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

can you find the derivative of the top and of the bottom?

OpenStudy (anonymous):

im so lost

OpenStudy (solomonzelman):

I would expand on the bottom first.

OpenStudy (solomonzelman):

how do you expand the \(\large\color{slate}{(4x+3)^2 }\) ?

OpenStudy (anonymous):

16x^2+24+9

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

24x you mean

OpenStudy (solomonzelman):

re-writing it, \(\large\color{slate}{-\left[ \displaystyle\lim_{x \rightarrow ~\infty}~\frac{(2x^2-1)}{16x^2+24x+9}\right] }\)

OpenStudy (solomonzelman):

find the derivative of top and bottom of the limit

OpenStudy (solomonzelman):

Do you know how to find the derivative of each of the terms in the denominator and the numerator?

OpenStudy (anonymous):

common number? is bottom 8(4x+3)????????

OpenStudy (solomonzelman):

you are trying to find the derivative?

OpenStudy (solomonzelman):

kay, lets start from here, what is the derivative of 2x^2 ?

OpenStudy (solomonzelman):

(Use the power rule)

OpenStudy (anonymous):

4x sorry it takes so long to reply I have to reconnect everytime

OpenStudy (solomonzelman):

yes, I understand, correct, what is the derivative of 1?

OpenStudy (anonymous):

1?

OpenStudy (solomonzelman):

Line y=3 will always have a slope of ?

OpenStudy (solomonzelman):

and any line of y=C, will have a slope of ?

OpenStudy (solomonzelman):



OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

So, what is derivative of -1, and of 9 ?

OpenStudy (anonymous):

0

OpenStudy (solomonzelman):

yes,

OpenStudy (solomonzelman):

i got to go in 10 minutes

OpenStudy (anonymous):

on no!!! can you help me finish this please, can u list steps so I can learn form it..please

OpenStudy (anonymous):

hello

OpenStudy (zale101):

@ingallina , Solomon was using L\hopital's rule, a simple way of solving limits, and since your function is in fractions, that's the best way to use l'hoptal's rule. If the limit is \(\large \frac{0}{0}\) and \(\large \frac{oo}{oo}\), these are the indeterminate forms. If you come up with an indeterminate answer for the limit, you then take the derivative of the top and bottom, if the result is still in indeterminate form, then take the second derivative, till you no longer have any indeterminate forms. L'Hopital's rule states: If the limit is \(\large \frac{0}{0}\) and \(\large \frac{oo}{oo}\) then \(\large\lim_{x \rightarrow a}\frac{f(x)}{g(x)}\frac{\large\bf Use~LHOPITAL'S~rule}{=}\lim_{x \rightarrow a}\frac{f'(x)}{g'(x)}\)

OpenStudy (anonymous):

Divide numerator and denominator by highest power term. In this case it is \(x^2\).